# Twilight: Adaptive Attention Sparsity with Hierarchical Top-$p$ Pruning

### v3

@@ 5.1 Accuracy Evaluation
Benchmarks and Models. We evaluate Twilight on two types of benchmarks: long-context, which includes Longbench bai2024longbench and RULER hsieh2024ruler , and medium-context (500 to 2k tokens), which includes GSM8K cobbe2021training , COQA reddy2019coqa , and the perplexity on the PG-19 dataset rae2019compressive . We select three widely used models, Longchat-7B-v1.5-32k li2023longchat , LLaMA2-7B-Chat touvron2023llama , and LLaMA-3.1-8B-Instruct dubey2024llama (128k context length), with two of them having long context ability $\geq$ 32k. They cover two mainstream attention implementations of multi-head attention (MHA) and group query attention (GQA) dubey2024llama .
Baselines. We use two SOTA top-$k$ sparse attention methods, Quest tang2024quest and DS yang2024post , and one SOTA non-top-$k$ method, MagicPIG chen2024magicpig , as our baselines. Following the baselines, we do not apply any sparse methods to the first two layers to ensure fair comparison. For DS, we use the optimized configurations tuned for each model provided by its official repository. The hyperparameter $p$ of Twilight is set to $0.95$ for LLaMA-2/3 and $0.85$ for Longchat, which will be explored in Section 5.3. Note that MagicPIG does not employ the budget mechanism but instead relies on two configurable parameters, $K$ and $L$, which directly influence its accuracy. In our experiments, we adopt two standard configurations from the original MagicPIG paper. Due to the lack of official MagicPIG support for LLaMA-2, we exclude these experiments from our evaluation.
@@
Results on Medium-Context Tasks. We then demonstrate that the Twilight Pruner itself does not negatively impact performance on two zero-shot generation tasks, GSM8K and COQA using the lm-harness framework gao2021framework , as well as one perplexity test on the PG-19 dataset. Since we are specifically evaluating the Pruner, we do not integrate Twilight into the baseline models. All the baselines use a budget of 128, which is comparable to the budget after Twilight’s pruning. The results in Table 4 show that Twilight outperforms Quest and DS by significant margins, with nearly zero loss compared to full attention.
@@ 5.2 Efficiency Evaluation
Datasets. We evaluate the efficiency of Twilight on both the self-attention operator and the end-to-end decoding stage on a single A100 GPU. We use Longbench, from which we select three different types of tasks: Qasper dasigi2021dataset for QA, GovReport huang2021efficient for summarization, and LCC guo2023longcoder for coding. We use prompts ranging from 10k to 30k tokens for evaluation. Given that Twilight is designed for deploying sparse attention in LLM serving systems, we use batch inference in our experiments.
Baselines and Implementation Details. We compare our methods with the following baselines: PyTorch’s scaled-dot-product-attention (SDPA), with FlashAttention2 (FA2) dao2023flashattention2 and Memory-Efficient Attention benjamin2022xformers as the backends; FlashInfer ye2025flashinfer , a high-performance kernel library for LLM serving; Quest, which achieves SOTA runtime performance among sparse attention methods. We integrate Twilight with both FlashInfer and Quest, resulting in FlashInfer-Twi and Quest-Twi. We modify the Quest kernels to support batch inference. We implement Twilight using both CUDA and OpenAI Triton tillet2019triton , following the technical details described in Section 4.2.

### v5

@@ 5.1 Accuracy Evaluation
Benchmarks and Models. We evaluate Twilight on two types of benchmarks: long-context, which includes Longbench [1] and RULER [16], and medium-context (500 to 2k tokens), which includes GSM8K [13], COQA [14], and the perplexity on the PG-19 dataset [15]. We select three widely used models, Longchat-7B-v1.5-32k [52], LLaMA2-7B-Chat [53], and LLaMA-3.1-8B-Instruct [54] (128k context length), with two of them having long context ability $\geq$ 32k. They cover two mainstream attention implementations of multi-head attention (MHA) and group query attention (GQA) [54].
Baselines. We use two SOTA top-$k$ sparse attention methods, Quest [9] and DS [12], and one SOTA non-top-$k$ method, MagicPIG [30], as our baselines. Following the baselines, we do not apply any sparse methods to the first two layers to ensure fair comparison. For DS, we use the optimized configurations tuned for each model provided by its official repository. The hyperparameter $p$ of Twilight is set to $0.95$ for LLaMA-2/3 and $0.85$ for Longchat, which will be explored in Section 5.3. Note that MagicPIG does not employ the budget mechanism but instead relies on two configurable parameters, $K$ and $L$, which directly influence its accuracy. In our experiments, we adopt two standard configurations from the original MagicPIG paper. Due to the lack of official MagicPIG support for LLaMA-2, we exclude these experiments from our evaluation.
@@
Results on Medium-Context Tasks. We then demonstrate that the Twilight Pruner itself does not negatively impact performance on two zero-shot generation tasks, GSM8K and COQA using the lm-harness framework [55], as well as one perplexity test on the PG-19 dataset. Since we are specifically evaluating the Pruner, we do not integrate Twilight into the baseline models. All the baselines use a budget of 128, which is comparable to the budget after Twilight’s pruning. The results in Table 4 show that Twilight outperforms Quest and DS by significant margins, with nearly zero loss compared to full attention.
@@ 5.2 Efficiency Evaluation
Datasets. We evaluate the efficiency of Twilight on both the self-attention operator and the end-to-end decoding stage on a single A100 GPU. We use Longbench, from which we select three different types of tasks: Qasper [56] for QA, GovReport [57] for summarization, and LCC [58] for coding. We use prompts ranging from 10k to 30k tokens for evaluation. Given that Twilight is designed for deploying sparse attention in LLM serving systems, we use batch inference in our experiments.
Baselines and Implementation Details. We compare our methods with the following baselines: PyTorch’s scaled-dot-product-attention (SDPA), with FlashAttention2 (FA2) [39] and Memory-Efficient Attention [59] as the backends; FlashInfer [47], a high-performance kernel library for LLM serving; Quest, which achieves SOTA runtime performance among sparse attention methods. We integrate Twilight with both FlashInfer and Quest, resulting in FlashInfer-Twi and Quest-Twi. We modify the Quest kernels to support batch inference. We implement Twilight using both CUDA and OpenAI Triton [60], following the technical details described in Section 4.2.